OpenStudy (anonymous):

Solve the system by substitution. y = 2x 3x + 2y = 21 A. x = 3, y = 6 B. x = 6, y = 3 C. x = 3, y = –6 D. x = –6, y = 3

OpenStudy (littlebird):

y=2x 3x+2(2x)=21 substitution 3x+4x=21 7x=21 x=3 Now use the first equation again y=2x y=2(3) y=6